The room numbers of two adjacent classrooms are two consecutive odd numbers. If their sum is 284, what are the classroom numbers?
Answer by oscargut(2103)

x+(x+2)=284 so 2x+2=284 then 2x=282 then x=141 and x+2=143
R: the numbers are 141 and 143